OpenStudy (kamille):

Hello, I need some help. I got the answer 9, but it should be 4,5. What am I doing wrong? (LOOK BELLOW!)

OpenStudy (anonymous):

I don't see anything here

OpenStudy (kamille):

Well, I got a=3, because V=a^3 then So right now we are looking at triangle ADC: \[AC ^{2}=3^{2}+3^2=18\] \[AC=\sqrt{18}=3\sqrt{2}\] then I need half of AC(lets call it AO) is \[AO=\frac{ 3\sqrt{2} }{ 2 }\] heighten of triange ADC (lets call it DO) is \[DO ^{2}=3^{2} - (\frac{ 3\sqrt{2} }{ 2 })^{2}=9-4.5=4.5\] \[DO=\sqrt{4,5}=\sqrt{\frac{ 9 }{2 }}=3\sqrt{\frac{ 1 }{2 }}\] So S of triangle ADC is S=DO*AC=9 V=(1/3)*9*3=9 BUT THE ANSWER SAYS IT SHOULD BE 4,5

OpenStudy (anonymous):

can you tell me the whole problem as well?

OpenStudy (kamille):

ahh, sorry! I need to find V of a pyramid!

OpenStudy (kamille):

butttt also, I have found my mistake S of a triangle is S=(ah)/2

OpenStudy (anonymous):

answer was 4,5? does that mean half of 9?

OpenStudy (kamille):

yup

OpenStudy (anonymous):

so you only need to substitute the S=ah with S=ah/2, which give you S=DO*AC/2=9/2

OpenStudy (anonymous):

do you get it?